# references.bib

### v1

@@ Abstract
To make sense of their surroundings, intelligent systems must transform complex sensory inputs to structured codes that are reduced to task-relevant information such as object category. Biological agents achieve this in a largely autonomous manner, presumably via self-super-vised learning. Whereas previous attempts to model the underlying mechanisms were largely discriminative in nature, there is ample evidence that the brain employs a generative model of the world. Here, we propose that eye movements, in combination with the focused nature of primate vision, constitute a generative, self-supervised task of predicting and revealing visual information. We construct a proof-of-principle model starting from the framework of masked image modeling (MIM), a common approach in deep representation learning. To do so, we analyze how core components of MIM such as masking technique and data augmentation influence the formation of category-specific representations. This allows us not only to better understand the principles behind MIM, but to then reassemble a MIM more in line with the focused nature of biological perception. From a theoretical angle, we find that MIM disentangles neurons in latent space, a property that has been suggested to structure visual representations in primates, without explicit regulation. Together with previous findings of invariance learning, this highlights an interesting connection of MIM to latent regularization approaches for self-supervised learning. The source code is available under https://github.com/RobinWeiler/FocusMIM
@@ 1 Introduction
Both biological and artificial intelligent systems must construct useful object representations (in the general case equivalent to classifiable) from large amounts of unlabeled data [lecun2022path]. In self-supervised learning (SSL), two main approaches have crystalized over the last years: multi-view based learning (via contrasting [oord2018representation], latent regularization [bardes2021vicreg] or distillation [grill2020bootstrap]), and masked image-modeling (MIM) [pathak2016context, he2022masked] which predicts occluded image content usually at the pixel-level, but to which also the more recently developed latent predictions [assran2023self] can be counted. While constrastive methods have been suggested as a model of SSL in the brain, structuring representational geometry via temporal [halvagal2023combination] and spatial cues [illing2021local], we postulate that the brain may be engaged in MIM via eye movements and attention shifts (Fig. 1): Perception, at a singular point in time, is selective, both through foveal vision and selective attention [simons1999gorillas]. But perception is also dynamic: saccadic eye movements and attention shifts flood the sensory stream with new information - in a manner that is to some degree predictable from knowledge about the direction and magnitude of the gaze shift (via a corollary discharge [crapse2008frontal, von1950reafferenzprinzip]) and image structure.
Figure 1: Representation learning through eye-movements. (a) Random patch masks used in artificial MIM approaches, gray patches are hidden from the network. (b) In our approach, selective masking is achieved through the inhomogeneous nature of foveal vision. Eye movements depicted as vectors reveal previously inaccessible or distorted information that is compared to the prediction $x^{t+\delta}$. (c) After pretraining on the prediction task, the quality of latent representations in the latent representation is assessed through classification accuracy in linear probing.
Studying the principles behind MIM is attractive as it requires less assumptions about temporal sequences and data augmentations than the mentioned multi-view approaches to biological SSL. Furthermore, such an approach fits well with predictive theories of cortical processing [dayan1995helmholtz, rao1999predictive, friston2005theory]: both predictive processing [pennartz2015brain] and MIM [kong2023understanding] conceptualize perception as the process of identifying a hierarchical latent variable model that could have caused the sensory inputs (this generative aspect is lacking in discriminative SSL models) and are predictive of future inputs. Experimentally, the frontal eye fields have been proposed to stabilize perception through predictions across eye movements [crapse2008frontal]. We postulate a larger role of this mechanism: the learning of visual features useful for downstream tasks such as classification.
@@ 2 Related Work
The breakthrough for MIM in representation learning arguably was achieved using random patch masks (Fig. 1a) in a vision transformer [he2022masked], an architecture that differs from information processing in biological networks. Subsequently, Kong et al. [kong2023understanding] and He et al. [he2022masked] found that an optimal learning ratio falls roughly between 40% and 80%, being sufficiently large but not overly aggressive.
A more biologically compatible CNN for MIM was developed by [pathak2016context]. The authors found that a central mask (predicting from peripheral context to a small central region) performed worse than randomly distributed patch or region masks, although quantitative reports were not provided. A more recent CNN-based MIM is ConvNext-V2 [woo2023convnext], which utilizes random patch masks like the transformer-based methods in combination with a shallow decoder. Xie et al. [xie2022simmim] further demonstrated the potential for MIM in CNNs, achieving high accuracy in a ResNet-based model. Neither Woo et al. [woo2023convnext] nor Xie et al. [xie2022simmim] investigated a classical self-supervised paradigm with limited amounts of labeled data and instead finetuned the pretrained model on the full ImageNet dataset used during pretraining.
Biological approaches to SSL derive mostly from multi-view approaches. They avoid representation collapse, a catastrophic network state in which different inputs are projected onto the same latent activity pattern and thus cannot be distinguished anymore, by introducing an expanding force that pushes representations apart. Illing et al. [illing2021local] used a single negative sample at a time, interpreted as being obtained by larger eye movements. Their approach converts contrastive predictive coding [oord2018representation] into a fully local learning rule. Interestingly, the model also has a generative component (predicting the next representation in a sequence of views) as a means to representation learning that is not investigated further. Local predictive learning (LPL) [halvagal2023combination], on the other hand relies on pushing representations apart based on their temporal distance, converting variance regularization [bardes2021vicreg] into a local learning rule. In contrast, MIM as we employ it, prevents representation collapse in a more minimalist way, i.e., without the necessity of negative examples [illing2021local] or accumulatie temporal statistics [halvagal2023combination], and instead solving a generative problem in which loss minimization is incompatible with a collapsed solution. As the approach presented here is an initial step, we use the spatially non-local backpropagation, however, see [lillicrap2020backpropagation, whittington2017approximation] and subsequent work for biologically plausible solutions yielding comparable weight updates.
@@ 3.1 Model
We reduce the task of predicting across eye-movements to its essence and avoid the burden of learning a spatial transformation, which we argue is orthogonal to the process of information restoration that drives representation learning [kong2023understanding]. As a consequence, inputs and predictions take place in the same reference frame (Fig. 1c), as in other MIM approaches [he2022masked, woo2023convnext, kong2023understanding]. The network architecture shown in Fig. 1c resembles an autoencoder where an encoder network maps input images to a compressed latent code from which a decoder network reconstructs the image. The encoder reduces the spatial dimension from 96x96 (with three color channels) to a spatial bottleneck of 12x12 (with 128 channels), forming the so called latent representation. Due to the large receptive fields of the latent space neurons at the output of the encoder, they map roughly onto neurons in visual area V4 or higher in primates [smith2001estimating]. From here, the inversely structured decoder (with deconvolutions instead of convolutions) maps the latent representations back onto a 96x96x3 output that constitutes the model’s expectation of image content. Both encoder and decoder consist of nine ResNet18-style blocks [he2016deep] in three stages without batch normalization layers, as convolutional networks are closer to biological networks than transformers.
@@ 3.2 Self-Supervised Pretraining
As other work in MIM, we define pretraining as the self-supervised training phase in which the model is presented with a partial view of an image and attempts to predict the image pixel by pixel in full resolution. The loss function is given as the mean squared reconstruction error in parts of visual space with incomplete inputs (masked or blurred, depending on the paradigm), normalized by the area over which the loss is calculated. This reconstruction loss is reminiscent of generative model learning in predictive coding [rao1999predictive]. Pretraining was conducted on the unlabeled split of the STL-10 dataset [coates2011analysis], which was specifically constructed for SSL. We investigated different masking strategies illustrated in Fig. 2 as inputs during pretraining:
@@
Random patches: The most common masking strategy used, e.g., in [he2016deep, woo2023convnext, kong2023understanding] (see also Fig. 1a). In our implementation, we split the image into 144 patches of 8x8 pixels and replaced the masked image content with the average color of the image.
Masked periphery: A central circle in full resolution with the periphery filled in the sample’s average color (see also Fig. 1b). Biologically, this strict discarding of out-of-focus information is imaginable through inattentiveness [simons1999gorillas].
@@
Figure 2: The masking strategies define the pretraining task. Top row: Exemplary inputs for different pretraining paradigms with masked areas covered in the image’s gray average color. Masked periphery and random patches are shown with 60% masking ratio. Middle row: Predictions from the given inputs after pretraining. Bottom image: Ground truth.
@@
Blurry periphery and Blurry random patches: Variations of the above paradigms with Gaussian-blurred information instead of average colors.
@@
Fovea filter: The circular center (”fovea”) of the image is in full resolution with a very small radius. The further out from the center a pixel is, the higher the level of blur, mimicking the physiological constraints of human foveal vision, i.e., the distribution of cone photoreceptors across the retina [purves2004neuroscience, p. 244]. We follow the implementation of [perry2002gaze, jiang2015salicon] by means of a resolution map determining the blur filter’s spatial frequency at a given distance from the center.
@@
For both random patches and masked periphery (using a square mask instead of a circle), we experimented with sparse convolutions as in [woo2023convnext] to avoid processing of borders between masks and image content. However, we found no significant difference in linear probing accuracy (as described below), so we continued with standard convolutions. During pretraining, we employed random-resized cropping (scale $\in[0.08,1.0]$, aspect ratio $\in[0.75,1.33]$). We used the ADAM optimizer [kingma2014adam] (weight decay=1e-8, batch size=512) and selected the learning rate from [5e-5, 1e-4, 5e-4] that led to the best classification performance (described below) for each of the major pretraining tasks (masked periphery, random patches, foveal filter). This value was then kept when blurry information was introduced. After 500 epochs of pretraining, we continued with the model achieving the lowest reconstruction loss for further analysis.
3.3 Linear Probing
@@
To analyze the quality of the learned representations after pretraining, linear probing was conducted. Here, we followed the standard procedure common in SSL of discarding the decoder and appending a linear layer to the frozen encoder. The 128x12x12x10 (latent dimension * classes of the STL-10 labeled splits) parameters of the linear layer were then fit without further augmentations on the training split of the STL-10 dataset (5,000 images), of which we used 1,000 for training validation and early stopping. We used images without mask or blurring, as these consistently led to higher readout accuracy. As in pretraining, we employed the ADAM optimizer [kingma2014adam] (learning rate=1e-4, weight decay=1e-8, batch size=512). Classification accuracy was then evaluated on the 8,000 test images, averaged across five randomly seeded runs. We also experimented with multilayer perceptron readout and finetuning (retraining the whole network), but neither of these approaches led to large differences in accuracy.
@@ 3.4 Correlation of Neurons in Latent Space
with $R$ defined as a tensor containing the $n$ latent representation vectors $\mathbf{r}_{i}$ (that are z-scored across the batch dimension), with $n$ equal to the number of samples, in our case $n=100.000$. As in [bardes2021vicreg], the sum of the squared off-diagonal entries in $C$ then quantifies the degree to which neurons are correlated:
@@
Unlike in the pretraining paradigms outlined so far, eye movements do not reveal all previously unavailable information at once. Instead, they usually fall on salient points, often objects [hayhoe2005eye, yarbus1967eye]. Objects may also be segmented already before recognition, either from motion signals [brucklacher2023learning], occlusion geometry, or both [tsao2022topological], making them the prioritized prediction targets. We thus evaluated the influence of reduced learning signal availability from the background. First, we obtained a figure-background segmentation mask using the rembg-library (https://github.com/danielgatis/rembg). We discarded images in which less than 100 pixels (out of 96x96) received a confidence score of 0.8 (out of 1) or higher. In these $\sim$8,000 images, no clear segmentation mask was obtained. Otherwise, as shown in Fig. 3, the masks were generally high-quality and only small object parts were missed. During pretraining, the reconstruction loss was then multiplied with the mask value (between 0 and 1) at each point in space and normalized to account for the reduced prediction area.
Figure 3: Representative presegmentation masks used to investigate how discarding loss signals from the background affects network performance, obtained with rembg. Displayed are the masks obtained with the rembg-library, with brightness indicating confidence, overlayed onto the respective STL-10 images.
3.6 Determining the Optimal Masking Ratio
For each of the major masking paradigms, we varied the masking ratio to choose a model with optimal classification accuracy. For the random patches, we found a masking ratio of 60% to lead to the best representations. For the masked periphery, a higher ratio of 80% performed slightly better. We thus continued with this fraction for the remaining results and refer to [kong2023understanding, he2016deep] for more detailed studies on the influence of the masking ratio.

### v2

@@ Abstract
To make sense of their surroundings, intelligent systems must transform complex sensory inputs to structured codes that are reduced to task-relevant information such as object category. Biological agents achieve this in a largely autonomous manner, presumably via self-super-vised learning. Whereas previous attempts to model the underlying mechanisms were largely discriminative in nature, there is ample evidence that the brain employs a generative model of the world. Here, we propose that eye movements, in combination with the focused nature of primate vision, constitute a generative, self-supervised task of predicting and revealing visual information. We construct a proof-of-principle model starting from the framework of masked image modeling (MIM), a common approach in deep representation learning. To do so, we analyze how core components of MIM such as masking technique and data augmentation influence the formation of category-specific representations. This allows us not only to better understand the principles behind MIM, but to then reassemble a MIM more in line with the focused nature of biological perception. We find that MIM disentangles neurons in latent space without explicit regularization, a property that has been suggested to structure visual representations in primates. Together with previous findings of invariance learning, this highlights an interesting connection of MIM to latent regularization approaches for self-supervised learning. The source code is available under https://github.com/RobinWeiler/FocusMIM
@@ 1 Introduction
Both biological and artificial intelligent systems must construct useful object representations (in the general case equivalent to classifiable) from large amounts of unlabeled data [lecun2022path]. In self-supervised learning (SSL), two main approaches have crystalized over the last years: multi-view based learning (via contrasting [oord2018representation], latent regularization [bardes2021vicreg] or distillation [grill2020bootstrap]), and masked image-modeling (MIM) [pathak2016context, he2022masked] which predicts occluded image content usually at the pixel-level, but to which also the more recently developed latent predictions [assran2023self] can be counted. While constrastive methods have been suggested as a model of SSL in the brain, structuring representational geometry via temporal [halvagal2023combination] and spatial cues [illing2021local], we postulate that the brain may be engaged in MIM via eye movements and attention shifts (Fig. 1): Perception, at a singular point in time, is selective, both through foveal vision and selective attention [simons1999gorillas]. Sequentially, saccadic eye movements and attention shifts then flood the sensory stream with new information - in a manner that is to some degree predictable from knowledge about the direction and magnitude of the gaze shift (via a corollary discharge [crapse2008frontal, von1950reafferenzprinzip]) and image structure.
Studying the principles behind MIM is attractive, as it requires fewer assumptions about temporal sequences and data augmentations than the mentioned multi-view approaches to biological SSL and achieves performance on par or close to contrastive methods [woo2023convnext]. MIM fits well with predictive theories of cortical processing [dayan1995helmholtz, rao1999predictive, friston2005theory]. Both conceptualize perception as the process of iden-
Figure 1: Representation learning through eye movements. (a) Random patch masks used in artificial MIM approaches, gray patches are hidden from the network. (b) In our approach, selective masking is achieved through the inhomogeneous nature of foveal vision. Eye movements depicted as vectors reveal previously inaccessible or distorted information that is compared to the prediction $x^{t+\delta}$. (c) After pretraining on the prediction task, the quality of latent representations in the latent representation is assessed through classification accuracy in linear probing.
tifying a hierarchical latent variable model that could have caused the sensory inputs [pennartz2015brain, kong2023understanding] (this generative aspect is lacking in discriminative SSL models) and are predictive of future inputs. Experimental evidence for predictions of visual content across saccades comes, e.g., from [ehinger2015predictions]. Anatomically, the frontal eye fields have been proposed to predict across eye movements to stabilize perception [crapse2008frontal]. We postulate a larger role of these predictions: the learning of visual features useful for downstream tasks such as classification.
@@ 2 Related Work
The breakthrough for MIM in representation learning arguably was achieved using random patch masks (Fig. 1a) in a vision transformer [he2022masked], an architecture that differs from information processing in biological networks. Investigations into the influence of masking strategies found that the optimal masking ratio falls roughly between 40% and 80%, being sufficiently large but not overly aggressive [kong2023understanding, he2022masked]. A more biologically compatible architecture (CNNs) for MIM was used by Pathak et al. [pathak2016context], as do we for this reason. Pathak et al. found that a central mask (predicting from peripheral context to a small central region) performed worse than randomly distributed patch or region masks, although quantitative reports were not provided. Here, we address this shortcoming and systematically study the choice of the masking paradigm. A recent CNN-based MIM is ConvNeXt-v2 [woo2023convnext], which utilizes random patch masks like the transformer-based methods in combination with a shallow decoder, whereas we strive to make the masking paradigm more plausible. Xie et al. [xie2022simmim] further demonstrated the potential for MIM in CNNs, achieving high accuracy in a ResNet-based model. In contrast to their evaluation paradigm that employs the same (full) dataset both during pretraining and fine-tuning [[, see also ]]woo2023convnext, we return to the ethologically relevant setting with limited amounts of labeled data.
Biological approaches to SSL derive mostly from multi-view approaches. They avoid representation collapse, a catastrophic network state in which different inputs are projected onto the same latent activity pattern and thus cannot be distinguished anymore, by introducing an expanding force that pushes representations apart. Illing et al. [illing2021local] used a single negative sample at a time, interpreted as being obtained by larger eye movements. Their approach converts contrastive predictive coding [oord2018representation] into a fully local learning rule. Interestingly, the model also has a generative component (predicting the next representation in a sequence of views) as a means to representation learning that is not investigated further. Local predictive learning (LPL) [halvagal2023combination], on the other hand, relies on pushing representations apart based on their temporal distance, converting variance regularization [bardes2021vicreg] into a local learning rule. In contrast, MIM, as we employ it, prevents representation collapse in a more minimalist way, i.e., without the necessity of negative examples [illing2021local] or accumulated temporal statistics [halvagal2023combination], but instead solves a generative problem in which loss minimization is incompatible with a collapsed solution. As the approach presented here is an initial step, we use the spatially non-local backpropagation, however, see [lillicrap2020backpropagation, whittington2017approximation] and subsequent work for biologically plausible solutions yielding comparable weight updates.
@@ 3.1 Model
We reduce the task of predicting across eye movements to its essence and avoid learning a remapping of spatial receptive fields, which we argue is orthogonal to the process of information restoration that drives representation learning [kong2023understanding]. Thus, inputs and predictions take place in the same reference frame (Fig. 1c), as in other MIM approaches [he2022masked, woo2023convnext, kong2023understanding]. The network architecture, shown in Fig. 1c, resembles an autoencoder. Unless explicitly noted otherwise, we use the following setup: The encoder reduces the spatial dimension from 96 x 96 (with three color channels) to 12 x 12 (x 128 channels). Due to the large receptive fields of the latent space neurons at the output of the encoder, they map roughly onto neurons in visual area V4 or higher in primates [smith2001estimating]. The inversely structured decoder (with deconvolutions instead of convolutions) maps the latent representations back onto a 96 x 96 x 3 output, constituting the model’s expectation of image content. Both encoder and decoder comprise nine ResNet18-style blocks [he2016deep] of two convolutions passed by a skip connection. Convolutions are followed by a ReLU activation function except the last decoder layer, which is followed by a sigmoid function. The convolution at the entry to each third block has stride = 2, which is where we double and half the number of channels in the encoder and decoder, respectively, starting with 32 channels at the first encoder block.
@@ 3.2 Self-Supervised Pretraining
As other work in MIM, we define pretraining as the self-supervised training phase in which the model is presented with a partial view of an image and attempts to predict the image pixel by pixel in full resolution. The loss function is given as the mean squared reconstruction error in parts of visual space with incomplete inputs (masked or blurred, depending on the paradigm), normalized by the area over which the loss is calculated. This generative loss is reminiscent of predictive coding [rao1999predictive]. Pretraining was conducted on the unlabeled split of the STL-10 dataset [coates2011analysis], which was specifically constructed for SSL with few labeled examples, and is sufficiently naturalistic and large to exclude the need for another dataset. We investigated a broad spectrum of input masking strategies (Fig. 2) that pertain to the original hypothesis of peripheral masking:
@@
Random patches: The most common masking strategy, used, e.g., in [he2016deep, woo2023convnext, kong2023understanding] (cf. Fig. 1a) serves as an optimal baseline. We split the image into 144 patches of 8 x 8 pixels and replaced the masked image content with the average color of the image.
@@
Figure 2: The masking strategies define the pretraining task. Top row: Exemplary inputs for different pretraining paradigms, with masked areas covered in the image’s gray average color. Masked periphery and random patches are shown with 60% masking ratio. Bottom row: Predictions from the given inputs after pretraining. Right image: Ground truth
Masked periphery: A central circle in full resolution, with the periphery masked with the sample’s average color (see also Fig. 1b). Biologically, this strict discarding of out-of-focus information is imaginable through inattentiveness [simons1999gorillas].
@@
Blurry periphery and Blurry random patches: Variations of the above paradigms with local Gaussian blurring instead of a uniformly colored mask to test the influence of incomplete masking.
@@
Foveal filter: A centered circle with a small radius in full resolution. Outside the circle, the level of blur increases with distance to the center, mimicking the physiological constraints of human foveal vision, i.e., the distribution of cone photoreceptors across the retina [purves2004neuroscience, p. 244]. The implementation follows [perry2002gaze, jiang2015salicon].
@@
For both random patches and masked periphery (using a square mask instead of a circle), we experimented with sparse convolutions as in [woo2023convnext] to avoid processing of borders between masks and image content. However, we found no significant difference in linear probing accuracy (as described below), so we continued with standard convolutions.
3.3 Choice of Pretraining Hyperparameters
During pretraining, we employed random-resized cropping (scale $\in$$\{0.08,1.0\}$, aspect ratio $\in$$\{0.75,1.33\}$). We used the Adam optimizer [kingma2014adam] with weight decay = 1e88-8- 8, batch size = 512, and selected the learning rate from {5e-5, 1e-4, 5e-4} that led to the best classification performance (described below) for each of the major pretraining tasks (masked periphery, random patches, foveal filter) to render the obtained findings independent of the choice. This value was then kept when blurry information was introduced. For each of the major masking paradigms, we additionally varied the masking ratio from 10 - 90% in steps of 10 to choose a model with optimal classification accuracy. For the random patches, we found a masking ratio of 60% to lead to the best representations. For the masked periphery, a higher ratio of 80% performed slightly better. We thus continued with this fraction for the remaining results, and refer to [kong2023understanding, he2016deep] for more detailed studies on the influence of the masking ratio. After 500 epochs of pretraining, we continued with the model that achieved the lowest reconstruction loss for further analysis.
3.4 Linear Probing to Assess Representation Quality
@@
To analyze the quality of the learned representations after pretraining, linear probing was conducted. Here, we followed the standard SSL procedure of discarding the decoder and appending a linear layer to the frozen encoder. The 128 x 12 x 12 x 10 (latent dimension * classes of the STL-10 labeled splits) parameters of the linear layer were then fit without further augmentations on the training split of the STL-10 dataset (5,000 images), of which we used 1,000 for training validation and early stopping at the minimum validation loss. Irrespective of pretraining paradigm, we used images without mask or blurring in this phase, as these consistently led to higher readout accuracy. As in pretraining, we employed the Adam optimizer [kingma2014adam] (learning rate = 1e-4, weight decay = 1e-8, batch size = 512). Classification accuracy was then evaluated on the 8,000 test images, averaged across five randomly seeded runs. We also experimented with multilayer perceptron readout and fine-tuning (retraining the whole network), but neither of these approaches led to large differences in accuracy.
3.5 Generalization to a different network architecture
To test in how far the influence of masking strategy depends on the network architecture, we pretrained a modern CNN, the ConvNeXt-v2-Pico [woo2023convnext]. Key differences to the ResNet described above are increased receptive field size, separate depthwise and pointwise convolutions, the use of GELU activation functions, and a shallow decoder consisting of a single block. Pretraining on the STL-10 dataset was conducted in two configurations: with random patch masks or with masked periphery. Due to the ConvNeXt-v2’s initial stage that splits the input image into 7 x 7 patches, we implemented peripheral masking as predicting from the central 3 x 3 patches to the surround. For the random patches, we used a masking ratio of 60%, as suggested in [woo2023convnext]. Models were then pretrained for 370 epochs with hyperparameters close to [woo2023convnext]: batch size = 512, learning rate selected for optimal performance out of {1.6e-4, 5e-4, 1.5e-3}, AdamW optimizer [loshchilov2017decoupled] with 20 epoch linear warmup and subsequent cosine decay, weight decay = 0.05 as well as random resized cropping and horizontal flipping augmentations. For linear probing and fine-tuning, random erasing was additionally applied, and warmup was reduced to three epochs, with a base learning rate of 0.0006 (fine-tuning) and 0.008 (linear probing). While fine-tuning led to the best performance when using global average pooling before the classification head, linear readout profited from skipping this, so we proceeded with the optimal settings for each.
@@ 3.6 Correlation of Neurons in Latent Space
with $R$ defined as a tensor containing the $n$ latent representation vectors $\mathbf{r}_{i}$ that are z-scored across the batch dimension, with $n$ equal to the number of samples (we compute across one mini-batch with $n=512$). As in [bardes2021vicreg], the sum of the squared off-diagonal entries in $C$ then quantifies the degree to which neurons are correlated:
@@
Figure 3: Representative presegmentation masks used to investigate how discarding loss signals from the background affects network performance. The masks were obtained with the rembg-library, with brightness indicating confidence, overlaid onto the respective STL-10 images.
@@
Unlike assumed in section 3.6, eye movements reveal information sequentially, often targeting salient points such as objects [hayhoe2005eye, yarbus1967eye]. Objects may also be segmented already before recognition, either from motion signals [brucklacher2023learning, tsao2022topological] or occlusion geometry [tsao2022topological], making them prioritized prediction targets. We thus evaluated the influence of reduced learning signal availability from the background. First, we obtained a figure-background segmentation mask using the rembg-library (https://github.com/danielgatis/rembg). We discarded $\sim$ 8,000 images for which no clear segmentation mask could be obtained. In these, less than 100 pixels (out of 96 x 96) received a confidence score of 0.8 (on a scale from 0 to 1) or higher. In the remaining images, the masks were generally high-quality and only small object parts were missed (Fig. 3). During pretraining, the reconstruction loss was then multiplied with the foreground confidence score at each point in space and normalized to compensate the reduced prediction area.